Find the equation to the straight line which are at a distance of 1unit from the origin and which pass through the point (3,1).

Grade:11

1 Answers

Arun
5 years ago
 
Dear student
 
let the line is 
y – 1 = m (x- 3)
 
y – mx + 3m – 1 = 0
 
Now equate this with perpendicular length 1 from orogin you will get the valuesof m